Blast Co. is considering the acquisition of a unit from the French government. Its initial outlay would be $3,900,000. It will reinvest all the earnings in the unit. It expects that at the end of 8 years, it will sell the unit for 13,700,000 euros after capital gains taxes are paid. The spot rate of the euro is $1.06 and is used as the forecast of the euro in the future years. Blast has no plans to hedge its exposure to exchange rate risk. The annualized U.S. risk-free interest rate is .05 regardless of the maturity of the debt, and the annualized risk-free interest rate on euros is .07, regardless of the maturity of debt. Assume that interest rate parity exists. Blast ’s cost of capital is 0.20. It plans to use cash to make the acquisition. Determine the NPV under these conditions.
Answer #1

Initial Outlay = $ 39,00,000

Amount Receive after 8 years = 1,37,00,000 euros

Spot rate: 1 euro = $ 1.06

Forward Rate using Interest rate parity theory

FR/SR = ( 1+ Rate Euro)8/ (1+ Rate $)8

FR/1.06 = (1.07)8/(1.05)8

FR = 1.2327

Amount Receive after 8 Years in Dollar = 13700000* 1.2327 = $ 1,68,88,157

NPV = Future Value/PVF(20%, 8 years) - Present Value

= 16888157 / PVF ( .20, 8) - 39,00,000

= 1,68,88,157/3.8372 - 39,00,000

= $ 5,01,213